# CLEAR: Towards Contextual LLM-Empowered Privacy Policy Analysis and Risk Generation for Large Language Model Applications

### v2

@@ 3.3. Data Analysis
All study sessions were audio-recorded and then transcribed for analysis. We followed established open coding procedures (Brod et al., 2009) to analyze the interview data. Two members of our research team independently coded 20% of the sample, generating a set of preliminary codes using MAXQDA. They compared their codes and reconciled any differences. Using the same codebook, the two researchers coded the rest of the data. In this process, they constantly compared and discussed their codes to ensure full agreement. As the coding process was discussion-based, inter-coder reliability was not necessary (McDonald et al., 2019). Using the final codebook, we conducted a thematic analysis to identify themes. The complete codebook is provided in Appendix.
@@ 3.4. Key Insights
Participants also questioned the purpose behind data collection. Some of them wondered if their data was used for improving services or for targeted advertising, which blurred the lines of user consent, as it was not clear to the users what they had consented for. Furthermore, another significant concern was the potential impacts on users’ privacy. A participant in G1 said “I really want to know if my personal health information will be shared with a credit card company or a life insurance company… I think that would be very helpful for users if they know such kind of information.
@@ 3.4.2. KI2: Expectations for Contextual, Simplified, and Focused Privacy Information
Participants expressed a desire for privacy information that is contextual, concise, and easy to understand, emphasizing the importance of clearly communicating the reasons for and consequences of data usage. For instance, participants in G4 suggest adding tooltips to show definitions of words in the policy (“if I am not really sure what this means, I can click on it to quickly get a definition for what that means.”), “visualizing important information on one page while providing links to additional details”, and using simple language to summarize the information concisely. Participants emphasized that they “don’t want to be exposed to all of the information”, instead, they expect “one page to include only the most important (privacy) information.” Furthermore, participants highlighted the need for privacy policies and risk alerts to be relevant to the context in which they are engaged. They proposed features like highlighting sensitive information, summarizing pertinent policy sections, and outlining potential privacy risks in a clear manner For example, one participant in G5 expected to have a tool to “highlight the key permissions you’re granting in bullet points which… you should be aware of at a glance, ensuring you know exactly what you’re consenting to.” Such brief and in-context privacy notices are also able to enhance participant understanding and trust towards LLMs.
@@ 3.5. Design Goals
  • DG1: Enhance Understanding of Data Access and Usage. It should help participants understand (1) what sensitive data is accessed by LLMs, (2) who accesses it, (3) why it is accessed, and (4) how it affects them. It should provide clear explanations of LLMs’ data sharing practices and its implications for the users (KI1 in Section 3.4.1).
  • DG2: Develop Contextual and User-Friendly Privacy Policies. It should develop privacy policies that not only succinct but also contextual and easy to comprehend. It should summarize important context-related information in the privacy policy into a single page that serves as a guide to help users access more detailed information if needed (KI2 in Section 3.4.2).
@@ 4.2. Example Use Case
In the second scenario (Fig. 3), another user, Randy would like to ask the Gemini plugin in Gmail to summarize an email. After identifying sensitive information in the email, CLEAR notifies Randy about the specific categories of sensitive data detected, providing related privacy policy excerpts and highlighting potential risks. The tool enables users to identify privacy risks, understand their implications, and make informed decisions in real-time while interacting with LLM-enabled applications.
@@
    Similarly, Randy accesses policy snippets and potential risks of the Gemini plugin by clicking ”Know more,” seeing who can access his data and evaluating potential risks, which aids his decision-making.
@@ 4.3.1. Identifying sensitive information
To ensure that users do not directly share their sensitive information with LLMs while using CLEAR, we first use Microsoft Presidio [^3] Microsoft Presidio: https://microsoft.github.io/presidio/ to identify the types of sensitive information in the user’s input. Microsoft Presidio runs locally without sending the information to the cloud. In the case of direct conversational interfaces for LLM (e.g., ChatGPT), the user’s input includes the prompt sent to the LLM. For LLM add-ons (e.g., the Gemini addon for Gmail), the user input also includes the user’s private data included in the context (e.g., the content of the underlying email). CLEAR passes only the types of information, rather than the actual information itself, to the LLM for the subsequent extraction of contextual privacy policies and the generation of potential privacy risks, as detailed below.
@@ 4.3.2. Extracting relevant privacy policy snippets
To achieve this, we utilize few-shot learning techniques (Brown et al., 2020), which enable our system to learn effectively from a limited set of examples and generalize this learning to identify relevant sections of privacy policies. This approach allows CLEAR to quickly adapt to different privacy policies and pinpoint the most crucial snippets. Specifically, these snippets detail who has access to the sensitive information and for what purposes, tailored to the types of information users have entered (KI1). This method not only speeds up the process but also enhances the contextual understanding for users, helping them better comprehend the potential privacy risks associated with inputting their data into LLMs.
Prompt for generating privacy policy snippets Given the privacy policy: {policy}, and the input sensitive information: {sensitive information}, cite the sentences in the privacy policy that are semantically related with the input sensitive information only during using the LLM service rather than creating an account. List who will access the data (service provider and third parties) and the reason for access. Each data accessor should be different. The language and vocabulary should be easy to understand for kids in elementary school.
@@ 4.3.3. Generating potential privacy risks
The final component of CLEAR involves generating potential privacy risks associated with sharing sensitive information with LLMs. We utilize a chain-of-thoughts approach (Wei et al., 2022), which involves sequentially reasoning through the potential implications of data sharing. This method is structured as follows:
  • Knowledge Extraction: CLEAR first uses LLMs to extract structured knowledge from academic papers that discuss LLM privacy risks. Table 3 shows a list of academic papers that we used for knowledge extraction. These papers were initially identified through an automated search process using Google Scholar, employing the keyword privacy risk AND (large language model OR LLM). After retrieving the relevant papers, we prioritized those that included surveys or user studies related to LLM and privacy risks. The process of knowledge extraction was semi-automated: LLMs were used to analyze the text of these papers, extracting key insights and structuring them into predefined categories such as privacy risks and their causes, influences, and evidence. Then, these structured outputs were manually verified by experts for accuracy and relevance. This combination of automated extraction and human validation contributes to the generalizability of the system by ensuring that the generated risks are both comprehensive and grounded in robust, well-researched sources. The structured knowledge generated from this process is detailed in Appendix, where examples of the extracted risks are provided.
@@
  • Risk Generation: Using the prompts, the LLM envisions and articulates potential privacy risks based on the context-of-use and the types of private data included in user inputs. This step involves outlining scenarios where sensitive information could be misused or inadvertently exposed, providing users with a comprehensive view of the risks involved. Building on the taxonomy of AI privacy risks proposed by Lee et al. (Lee et al., 2024), we tag each generated privacy risk with its corresponding risk category. Inspired by DG2, each privacy risk has a brief summary and an expandable detailed explanation. This method ensures users receive a clear, context-aware understanding of the privacy risks they face when interacting with LLMs.
@@ 5.2.1. Use case of ChatGPT
We employed a within-subjects study design. Each participant was asked to enter two prompts to ChatGPT, one before using CLEAR and one after using it, and then compared their experiences. The detailed procedures are presented below.
Each session contained the following parts. After the consent, we introduced the team and the study objectives. We then conducted a pre-intervention assessment. We asked the participants to enter one of the following two prompts in ChatGPT, then share their perceptions about who would access what information, why, and how it would affect them. We selected these two prompts because both represented a typical scenario of interacting with LLMs that involve user inputs containing personally identifiable information. To control for potential biases, we counterbalanced the order in which the participants reviewed the two prompts.
After the pre-intervention assessment, we introduced CLEAR to our participants and asked them to use it freely for five minutes to explore its feature and get familiar. Then, participants entered the other prompt to ChatGPT with CLEAR enabled. At this time, the CLEAR interface should be triggered as the participants entered the prompts and provide additional information as described in Section 4. We asked participants to explore the information provided by CLEAR. When they finished, we asked about their perceptions again about who would access what information, why, and how it would affect them.
After the two conditions, we interviewed the participants, asking them to compare their experiences with and without CLEAR and explain the differences, if any, and how the policy snippets and potential privacy risks impacted their understanding and potential actions.
Finally, participants completed a post-study questionnaire that included the System Usability Scale (SUS) and additional questions on their the overall user experience and the quality and consistency of the contents presented to them.
@@ 5.2.2. Use case of the Gemini plugin in Gmail
We randomly selected six emails from the Enron public email dataset [^11] https://www.cs.cmu.edu/ enron/ as the dataset for case study. Since LLM-generated synthetic data is widely used for data augmentation (Schmidhuber and Kruschwitz, 2024), and to ensure that the selected emails contained sensitive information, we used GPT-4 to synthesize the sensitive information and integrated it into the selected six emails.
Each session contained the following parts. After the consent, we began with introductions, where the research team and the objectives of the study were presented.
The second part was a pre-intervention assessment. We asked the participants to review and summarized the content of three emails randomly selected from the six emails in the dataset. Based on this experience, we asked participants’ perceptions about who would access what information, why, and how it would affect them. To control for potential biases, we counterbalanced the order in which the participants encounter the emails.
@@
Similar to the first case study, after the two conditions, we interviewed the participants, asking them to compare their experiences with and without CLEAR and explain the differences, if any, and how the policy snippets and potential privacy risks impacted their understanding and potential actions. Finally, participants completed a post-study questionnaire that included the System Usability Scale (SUS) and additional questions on their the overall user experience and the quality and consistency of the contents presented to them.
@@ 5.3. Data analysis
We followed the same approach for the data analysis for the results of both case studies. For the user rating data (i.e., SUS scores), we computed the mean and standard deviation for each statement. For the interview data, two coders independently reviewed and coded the transcripts, following the method described in Section 3.3. After completing their individual coding, the coders iteratively refined the codebook through discussion, identifying recurring themes. These themes included changes in privacy perceptions after using CLEAR, reactions to the prompts, and suggestions or feedback on the tool. The detailed codebooks used and refined during this analysis are available in Appendix. This structured approach to data analysis ensured that we could systematically capture and interpret both the quantitative and qualitative feedback on CLEAR, providing a comprehensive understanding of its impact on users’ privacy awareness and behavior.
@@ 6.1. System Usability
The evaluation of CLEAR in the context ChatGPT, based on the results of the System Usability Scale(SUS) questionnaire (Bangor et al., 2009), indicates a generally positive user experience, as shown in Fig 4. Participants rated each item on a 5-point Likert scale. They find the system easy to use and well-integrated, with most expecting that it would be quickly learned. The confidence in using the system is high, and the information provided is deemed useful. Additionally, policy snippets and privacy risks are considered appropriate and well-matched with the sensitive information. Participants did not perceive the system as complex, cumbersome, or annoying, nor did they feel a need for technical support.
@@ 6.2.1. Increase participants’ awareness of data practices in LLMs (P1, P2, P4, P6, P7, P9, P11, P12)
Meanwhile, participants also showed a deeper understanding of the reasons why their sensitive information is be accessed. For example, P9 initially thought that data breaches were primarily due to hackers: “I think the hackers will have access to this information because this is personal information that should not be put on a public website unless you are assured that your details are safe.” Later, they understood that many more entities could access their data, which amplified the risks for data breaches: “Vendors and service providers, OpenAI, customer service, training team, analytics team, affiliates are also gonna see your information. I think this is because these entities work on the ChatGPT platform. So they’re gonna see the information. The best way to curb this is to not share your personal information at all.” This newfound clarity extended to data access purposes, and participants became more aware that their information was used not only for service improvement, but also for training models and other purposes.
@@ 6.2.3. Lead to an intention for behavior change (P4, P6, P9)
The increased awareness and understanding fostered by the use of CLEAR led our participants to consider changing their behavior to better protect their privacy. P6, for example, concluded that using LLMs necessitates more cautious sharing of personal information: “The user should reconsider providing the name, address, and email address with the AI and just use the one-sentence summary that he/she is requesting help with…I need to take steps to make sure that I’m at least not providing over information about myself other than basic things for like updating a resume”. P9’s statement further highlights the importance of not sharing personal information publicly, “The government authorities will be able to access the data, so I think it is wise to not put your personal information on there.” This collective change in behavior reflects the participants’ intentions to adopt more privacy-conscious practices as a result of their improved understanding of LLM privacy risks.
@@ 6.3. Participants’ Reactions to the Disclosure of Sensitive Information
Our participants expressed a strong desire to remove sensitive information from their prompts or substitute it with fake data. P1 explicitly states their preference for removing all personal information: “I think I would remove all the all the personal information.” Similarly, P6 emphasizes the importance of excluding personal details such as name, email address, telephone number, and home address: “I just removed like my name my email address, my telephone number. And my home address from the prompt.” Additionally, P5 highlights an alternative approach by suggesting the use of fake contact details: “I may like disguise him by providing fake email and fake phone number. I will not provide these details and only give a generic name, email, and phone number.” These responses indicate a common practice among our participants to ensure their privacy and security by omitting or obfuscating sensitive data in their prompts. Some participants even suggested that CLEAR could directly provide solutions for them, as P4 mentioned, “I would like to know more like not only the information they present but also some solutions that I can choose.
However, several participants chose to proceed with sending the sensitive information without any modification. For instance, P13 stated, It wouldn’t stop me from using ChatGPT. I’ll send it out.” Because they believe that this information will still be obtained by other applications in other ways, regardness of whether it is not shared here. This opinion is further explained by P5: Because I already know that they [other online services] have access to this information. These days, your email and your phone number or your mailing address are some of the easiest things to get. So, I’m okay that if they are using so.”
@@
Participants showed a strong desire for CLEAR to identify more modalities and types of sensitive information. P2 expressed a need for the system to identify not only text but also sensitive information in images and PDF files: “I want it can not only identify this text but also the things in the image and PDF files.” P5 emphasized the importance of support for identifying sensitive information related to financial transactions and health issues: “there should be some support for [identifying] like financial transaction and health-related issues.” P1 raised concerns about the system’s ability to handle less directly sensitive information, questioning how the system would react to seemingly innocuous details like the names of siblings or past residences: “I just wonder with information that isn’t so directly sensitive. I wonder how the system would act. let’s say I input, I have three sisters and their names. Would that be a sort of red flag for your system? Or let’s say I lived in a city when I was younger. Is that a risk because I am not living there now? So I don’t know how the system would react. That’s what I mean by less sensitive.” These insights reflect participants’ interest in a more comprehensive and nuanced approach to identifying and managing sensitive information.

### v3

@@ 3.3. Data Analysis
All study sessions were audio-recorded and then transcribed for analysis. We followed established open coding procedures (Brod et al., 2009) to analyze the interview data. Two members of our research team independently coded 20% of the sample, generating a set of preliminary codes using MAXQDA. They compared their codes and reconciled any differences. Using the same codebook, the two researchers coded the rest of the data. In this process, they constantly compared and discussed their codes to ensure full agreement. As the coding process was discussion-based, inter-coder reliability was not necessary (McDonald et al., 2019). Using the final codebook, we conducted a thematic analysis to identify themes. The complete codebook is provided in the Appendix.
@@ 3.4. Key Insights
Participants also questioned the purpose of the data collection. Some of them wondered if their data was used for improving services or for targeted advertising, which blurred the lines of user consent, as it was not clear to users what they had consented to. Furthermore, another significant concern was the potential impacts on users’ privacy. A participant in G1 said “I really want to know if my personal health information will be shared with a credit card company or a life insurance company… I think that it would be very helpful for users if they had such information.
@@ 3.4.2. KI2: Expectations for Contextual, Simplified, and Focused Privacy Information
The participants expressed a desire for contextual, concise, and easy-to-understand privacy information, highlighting the importance of clearly communicating the reasons and consequences of data use. For instance, participants in G4 suggest adding tooltips to show definitions of words in the policy (“if I am not really sure what this means, I can click on it to quickly get a definition for what that means.”), “visualizing important information on one page while providing links to additional details”, and using simple language to summarize the information concisely. Participants emphasized that they “don’t want to be exposed to all the information”, instead they expect “one page to include only the most important (privacy) information.” Furthermore, participants highlighted the need for privacy policies and risk alerts to be relevant to the context in which they are engaged. They proposed features like highlighting sensitive information, summarizing pertinent policy sections, and outlining potential privacy risks in a clear manner For example, one participant in G5 expected to have a tool to “highlight the key permissions you are granting in bullet points which… you should be aware of at a glance, ensuring you know exactly what you’re consenting to.” Such brief and in-context privacy notices could also enhance participant understanding and trust towards LLMs.
@@ 3.5. Design Goals
  • DG1: Enhance Understanding of Data Access and Usage. It should help participants understand (1) what sensitive data is accessed by LLMs, (2) who accesses it, (3) why it is accessed, and (4) how it affects them. It should provide clear explanations of LLMs’ data sharing practices and their implications for the users (KI1 in Section 3.4.1).
  • DG2: Develop Contextual and User-Friendly Privacy Policies. It should develop privacy policies that are not only succinct but also contextual and easy to comprehend. It should summarize important context-related information in the privacy policy into a single page that serves as a guide to help users access more detailed information if needed (KI2 in Section 3.4.2).
@@ 4.2. Example Use Case
In the second scenario (Fig. 3), another user, Randy would like to ask the Gemini plugin in Gmail to summarize an email. After identifying sensitive information in the email, CLEAR notifies Randy about the specific categories of sensitive data detected, providing related privacy policy excerpts and highlighting potential risks. The tool enables users to identify privacy risks, understand their implications, and make informed decisions in real time while interacting with LLM-enabled applications.
@@
    Similarly, Randy accesses policy snippets and potential risks of the Gemini plugin by clicking “Know more,” seeing who can access his data and evaluating any potential risks, which aids his decision-making.
@@ 4.3.1. Identifying sensitive information
To ensure that users do not directly share their sensitive information with LLMs while using CLEAR, we first use Microsoft Presidio [^3] Microsoft Presidio: https://microsoft.github.io/presidio/ to identify the types of sensitive information in the user’s input. In the case of direct conversational interfaces for LLM (e.g., ChatGPT), the user’s input includes the prompt sent to the LLM. For LLM add-ons (e.g., the Gemini addon for Gmail), the user input also includes the user’s private data in the context (e.g., the content of the underlying email). CLEAR passes only the types of information, rather than the actual information itself, to the LLM for the subsequent extraction of contextual privacy policies and the generation of potential privacy risks, as detailed below.
@@ 4.3.2. Extracting relevant privacy policy snippets
To achieve this, we use few-shot learning techniques (Brown et al., 2020), which enable our system to learn effectively from a limited set of examples and generalize this learning to identify relevant sections of privacy policies. This approach allows CLEAR to quickly adapt to different privacy policies and pinpoint the most crucial snippets. Specifically, these snippets detail who has access to the sensitive information and for what purposes, tailored to the types of information users have entered (KI1). This method speeds up the process and enhances the contextual understanding for users, helping them better comprehend the potential privacy risks associated with entering their data into LLMs.
Prompt for generating privacy policy snippets Given the privacy policy: {policy}, and the input sensitive information: {sensitive information}, cite the sentences in the privacy policy that are semantically related to the input sensitive information only when using the LLM service rather than creating an account. List who will access the data (service provider and third parties) and the reason for access. Each data accessor should be different. The language and vocabulary should be easy to understand for kids in elementary school.
@@ 4.3.3. Generating potential privacy risks
The final component of CLEAR involves generating potential privacy risks associated with sharing sensitive information with LLMs. We use a chain-of-thoughts approach (Wei et al., 2022), which involves sequentially reasoning through the potential implications of data sharing. This method is structured as follows:
  • Knowledge Extraction: CLEAR first uses LLMs to extract structured knowledge from academic papers that discuss LLM privacy risks. Table 3 shows a list of academic papers that we used for knowledge extraction. These papers were initially identified through an automated search process using Google Scholar using the keywords privacy risk AND (large language model OR LLM). After retrieving the relevant papers, we prioritized those that included surveys or user studies related to LLM and privacy risks. The knowledge extraction process was semi-automated: LLMs were used to analyze the text of these papers, extracting key insights and structuring them into predefined categories such as privacy risks and their causes, influences, and evidence. Then, these structured outputs were manually verified by experts for accuracy and relevance. This combination of automated extraction and human validation contributes to the generalizability of the system by ensuring that the generated risks are both comprehensive and grounded in robust, well-researched sources. The structured knowledge generated from this process is detailed in the Appendix, where examples of the extracted risks are provided.
@@
  • Risk Generation: Using the prompts, the LLM envisions and articulates potential privacy risks based on the context of use and the types of private data included in user inputs. This step involves outlining scenarios where sensitive information could be misused or inadvertently exposed, providing users with a comprehensive view of the risks involved. Building on the taxonomy of AI privacy risks proposed by Lee et al. (Lee et al., 2024), we tag each generated privacy risk with its corresponding risk category. Inspired by DG2, each privacy risk has a brief summary and an expandable detailed explanation. This method ensures users receive a clear, context-aware understanding of the privacy risks they face when interacting with LLMs.
@@ 5.2.1. Use case of ChatGPT
We employed a within-subjects study design. Each participant was asked to enter two prompts to ChatGPT, one before using CLEAR and one after using it, and then compared their experiences. Below are the detailed procedures.
Each session contained the following parts. After obtaining consent, we introduced the team and study objectives, followed by a pre-intervention assessment. We asked the participants to enter one of the following two prompts in ChatGPT, and then share their perceptions about who would access what information, why, and how it would affect them. We selected these two prompts because both represented a typical scenario of interacting with LLMs that involve user inputs containing personally identifiable information. The prompts were counterbalanced to mitigate biases.
After the pre-intervention assessment, we introduced CLEAR to our participants and asked them to use it freely for five minutes to explore its features and get familiar. Then, participants entered the other prompt to ChatGPT with CLEAR enabled. At this time, the CLEAR interface was triggered as the participants entered the prompts and provided additional information as described in Section 4. Participants reviewed this information and shared their updated perceptions.
Following both conditions, we conducted interviews to compare their experiences with and without CLEAR, focusing on differences and the impact of policy snippets and privacy risk details on their understanding and actions.
Finally, participants completed a post-study questionnaire, including the System Usability Scale (SUS) and questions on overall experience, content quality, and consistency.
@@ 5.2.2. Use case of the Gemini plugin in Gmail
We randomly selected six emails from the Enron public email dataset [^11] https://www.cs.cmu.edu/~enron/ as the dataset for the case study. Since LLM-generated synthetic data is widely used for data augmentation (Schmidhuber and Kruschwitz, 2024), and to ensure that the selected emails contained sensitive information, we used GPT-4 to synthesize the sensitive information and integrated it into the selected six emails.
Each session contained the following parts. After the consent, we began with introductions, where the research team and the objectives of the study were presented. The second part was a pre-intervention assessment. We asked the participants to review and summarize the content of three emails randomly selected from the six emails in the dataset. Based on this experience, we asked participants’ perceptions about who would access what information, why, and how it would affect them. To control for potential biases, we counterbalanced the order in which the participants encountered the emails.
@@
Similar to the first case study, after the two conditions, we interviewed the participants, asking them to compare their experiences with and without CLEAR and explain the differences, if any, and how the policy snippets and potential privacy risks impacted their understanding and potential actions. Finally, participants completed a post-study questionnaire that included the System Usability Scale (SUS) and questions about their overall user experience and the quality and consistency of the contents presented to them.
@@ 5.3. Data analysis
We followed the same approach for the data analysis for the results of both case studies. For the user rating data (i.e., SUS scores), we computed the mean and standard deviation for each statement. For the interview data, two coders independently reviewed and coded the transcripts, following the method described in Section 3.3. After completing their individual coding, the coders iteratively refined the codebook through discussion, identifying recurring themes. These themes included changes in privacy perceptions after using CLEAR, reactions to the prompts, and suggestions or feedback on the tool. The detailed codebooks used and refined during this analysis are available in the Appendix. This structured data analysis approach ensured that we could systematically capture and interpret both the quantitative and qualitative feedback on CLEAR, providing a comprehensive understanding of its impact on users’ privacy awareness and behavior.
@@ 6.1. System Usability
The evaluation of CLEAR in the context of ChatGPT, based on the results of the System Usability Scale(SUS) questionnaire (Bangor et al., 2009), indicates a generally positive user experience, as shown in Fig 4. Participants rated each item on a 5-point Likert scale. They find the system easy to use and well-integrated, with most expecting that it would be quickly learned. The confidence in using the system is high, and the information provided is deemed useful. Additionally, policy snippets and privacy risks are considered appropriate and well-matched with the sensitive information. Participants did not perceive the system as complex, cumbersome, or annoying, nor did they feel a need for technical support.
@@ 6.2.1. Increase participants’ awareness of data practices in LLMs (P1, P2, P4, P6, P7, P9, P11, P12)
Meanwhile, participants also showed a deeper understanding of the reasons why their sensitive information is accessed. For example, P9 initially thought that data breaches were primarily due to hackers: “I think the hackers will have access to this information because this is personal information that should not be put on a public website unless you are assured that your details are safe.” Later, they understood that many more entities could access their data, which amplified the risks for data breaches: “Vendors and service providers, OpenAI, customer service, training team, analytics team, affiliates are also gonna see your information. I think this is because these entities work on the ChatGPT platform. So they’re gonna see the information. The best way to curb this is to not share your personal information at all.” This newfound clarity extended to data access purposes, and participants became more aware of their information being used not only for service improvement but also for training models and other purposes.
@@ 6.2.3. Lead to an intention for behavior change (P4, P6, P9)
The increased awareness and understanding fostered by the use of CLEAR led our participants to consider changing their behavior to better protect their privacy. P6, for example, concluded that using LLMs necessitates more cautious sharing of personal information: “The user should reconsider providing the name, address, and email address with the AI and just use the one-sentence summary that he/she is requesting help with…I need to take steps to make sure that I’m at least not providing over information about myself other than basic things like updating a resume”. P9’s statement further highlighted the importance of not sharing personal information publicly, “The government authorities will be able to access the data, so I think it is wise to not put your personal information on there.” This collective change in behavior reflects the participants’ intentions to adopt more privacy-conscious practices as a result of their improved understanding of LLM privacy risks.
@@ 6.3. Participants’ Reactions to the Disclosure of Sensitive Information
Our participants expressed a strong desire to remove sensitive information from their prompts or substitute it with fake data. P1 explicitly stated their preference for removing all personal information: “I think I would remove all the all the personal information.” Similarly, P6 emphasized the importance of excluding personal details such as name, email address, telephone number, and home address: “I just removed like my name my email address, my telephone number. And my home address from the prompt.” Additionally, P5 highlighted an alternative approach by suggesting the use of fake contact details: “I may like disguise him by providing fake email and fake phone number. I will not provide these details and only give a generic name, email, and phone number.” These responses indicate a common practice among our participants to ensure their privacy and security by omitting or obfuscating sensitive data in their prompts. Some participants even suggested that CLEAR could directly provide solutions for them, as P4 mentioned, “I would like to know more like not only the information they present but also some solutions that I can choose.
However, several participants chose to proceed with sending the sensitive information without any modification. For instance, P13 stated, It wouldn’t stop me from using ChatGPT. I’ll send it out.” They believed that this information could still be obtained by other applications in other ways, regardless of whether it was shared here or not. This opinion was further explained by P5: Because I already know that they [other online services] have access to this information. These days, your email and your phone number or your mailing address are some of the easiest things to get. So, I’m okay that if they are using so.”
@@
Participants showed a strong desire for CLEAR to identify more modalities and types of sensitive information. P2 expresses a need for the system to identify not only text but also sensitive information in images and PDF files: “I want it can not only identify this text but also the things in the image and PDF files.” P5 emphasizes the importance of support for identifying sensitive information related to financial transactions and health issues: “there should be some support for [identifying] like financial transaction and health-related issues.” P1 raised concerns about the system’s ability to handle less directly sensitive information, questioning how the system would react to seemingly innocuous details like the names of siblings or past residences: “I just wonder with information that isn’t so directly sensitive. I wonder how the system would act. Let’s say I input, I have three sisters and their names. Would that be a sort of red flag for your system? Or let’s say I lived in a city when I was younger. Is that a risk because I am not living there now? So I don’t know how the system would react. That’s what I mean by less sensitive.” These insights reflect participants’ interest in a more comprehensive and nuanced approach to identifying and managing sensitive information.